OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -3 2x - 6y = -6 A. -1 B. zero<-- I think C. 1 D. 2

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@GabeBae

OpenStudy (*ar*):

i think its B

OpenStudy (anonymous):

I plugged it in for x and it worked but it says y and nothing works if you plug it in for y.

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

@OwenOVOXO

OpenStudy (anonymous):

B was wrong.

OpenStudy (trojanpoem):

You can either multiply by inverse coefficients or get y and substitute 5x - 3y = -3 < * -6 2x - 6y = -6 * 3 -30x + 18 y = 18 6x - 18y = -18 -24x = 0 x = 0 -3y = -3 y = 1

OpenStudy (anonymous):

Thank you!!

OpenStudy (trojanpoem):

Any time.

OpenStudy (*ar*):

sorry i looked it up